At two points P and Q on a screen in Young`s double slit experiment, waves from slit S1 and S2 have a path difference of 0 and lambda/4 respectively. Then ratio of intensities at P and Q will be ?

Grade:12th pass

2 Answers

Sher Mohammad IIT Delhi
9 years ago
path difference 0,lambda/4
so phase difference 0, pi/2

in case of zero, intensity= (2*sqrt(I))^2=4I
in cese of pi/2 I+I+2*sqrt(I*I)cos(pi/2)=2I

so I at P/I at Q=2


thanks and regards
Sher MOhammad
faculty askiitians
Kushagra Madhukar
3 years ago
Dear student,
Pleae find the solution to your problem.
 
path difference = 0, λ/4
so phase difference = 0, π/2
 
in case of 0 phase difference, intensity = (2I)2 = 4I
in cese of π/2 phase difference, I + I + 2(I*I)cos(π/2) = 2I
so (I at P):(I at Q) = 2:1
 
Thanks and regards,
Kushagra